# Towards Serverless Processing of Spatiotemporal Big Data Queries

### v1

@@ Abstract.
Spatiotemporal data are being produced in continuously growing volumes by a variety of data sources and a variety of application fields rely on rapid analysis of such data. Existing systems such as PostGIS or MobilityDB usually build on relational database systems, thus, inheriting their scale-out characteristics. As a consequence, big spatiotemporal data scenarios still have limited support even though many query types can easily be parallelized.
In this paper, we propose our vision of a native serverless data processing approach for spatiotemporal data: We break down queries into small subqueries which then leverage the near-instant scaling of Function-as-a-Service platforms to execute them in parallel. With this, we partially solve the scalability needs of big spatiotemporal data processing.
@@ 1. Introduction
Aside from data values such as a temperature measurement or a boolean “something has a value here”, spatiotemporal data contain additional space and time attributes (Alam et al., 2022). Here, space attributes refers to locations usually indicated by latitude and longitude values and sometimes an additional geofence indicating a particular area for 2D spatial data or its 3D variant with an additional elevation level or 3D shapes. Time attributes, in contrast, provide a temporal context, usually based on timestamps but possibly also based on timespans. Such spatiotemporal data are usually encoded as either trajectory or raster data and stored in respective database systems. In this paper, we focus on trajectory data.
Nowadays, spatiotemporal data are being produced in exponentially growing volumes in a variety of application scenarios (Alam et al., 2022). Examples of spatiotemporal data include mobility traces of cars, cyclists, pedestrians, planes, or ships in transport, movement patterns of wildlife in animal observation, Earth observation data, or satellite movement data (Karakaya et al., 2020, 2022; Liang et al., 2024). While there are different ways for processing these data, e.g., in batch mode, many application domains rely on interactive analysis of big spatiotemporal data with queries that filter and aggregate by space and time attributes at the same time.
Existing solutions for this build on PostGIS or MobilityDB and, thus, inherit the limited scalability of the underlying relational PostgreSQL (Rese et al., 2025). Recent work by Bakli et al. (Bakli et al., 2025) here represents the state-of-the-art. Through complex sharding mechanisms and careful query planning, they managed to scale MobilityDB to up to eight machines – an impressive feat.
@@
At the same time, many query types can easily be broken down into vast numbers of subqueries operating on small data shards only, e.g., based on geographic regions or other spatial or temporal attributes, thus, parallelizing the query execution. Today, the go-to solution for these “embarrassingly parallel” (Hellerstein et al., 2019) tasks is Function as a Service (FaaS) in which developers write simple, stateless functions whose operation including on-demand scaling is then fully managed by the cloud provider (Bermbach et al., 2020, 2021).
In this paper, we hence sketch out our vision of a serverless data processing approach for spatiotemporal big data analytics, thus, leverage the rapid on-demand scaling of FaaS function instances for parallelizing query processing. For this, we propose to break down queries into subqueries following the MapReduce model(Dean and Ghemawat, 2008), automatically deploy these on a FaaS platform, and merge the set of subresults into a single response (see Figure 1). In our proposed approach, a dedicated coordinator function analyzes the query and then breaks it down based on spatial or temporal characteristics or a combination of both.
This idea is not completely new, in fact, our work was partially inspired by similar approaches for serverless processing of tabular data (Lounissi et al., 2025; Bodner et al., 2025). To our knowledge, it has not yet been proposed for spatiotemporal data – a data type which needs dedicated sharding strategies which can consider spatial and temporal characteristics. Existing sharding approaches are not directly applicable. The reason for this is that storing spatial data already means reducing 2D or 3D data to a 1D sequence of values, which also makes indexing hard (Rese et al., 2025), while the time context adds at least one more dimension.
Processing spatiotemporal big data on a single node, however, leads to slow response times as it stores and processes all data alone leading to nodes becoming slow which impacts any other workloads.
@@ 2. Approach
We propose to use a shared storage as data layer and shift query processing to FaaS applications. Striving for a modular design which allows for later integration of our approach into diverse architectures, i.e., independent of the implemented execution framework, we use FaaS, blob storage. Our architecture design, depicted in Figure 2, is as follows.
A spatiotemporal query initiated by the client, e.g., “Identify the flight-heaviest regions in the last six months” is our starting point. A starter function provides an HTTP endpoint to receive the query and invokes the centralized coordinator. The coordinator splits the query into subqueries and initializes the worker functions. Queries such as SELECT can easily be distributed and executed on data shards, while others, e.g., ORDER BY or AVG, depend on context, i.e., they require the complete result of the previous subquery to return correct result and therefore can only be executed by single node. This is why we design our architecture in a way that the can support all kinds of queries, parallelizable and non-parallelizable. For that, the coordinator manages waves of worker functions for each subquery with the amount of worker functions per wave dependent on the parallelizability of the subquery.
@@ 3. Conclusion & Future Work
In this paper, we proposed our vision of a native serverless data processing approach for spatiotemporal data: The proposed approach breaks down queries into small subqueries which then leverage the near-instant scaling of Function-as-a-Service platforms for parallel execution.

### v2

@@ Abstract.
Spatiotemporal data are being produced in continuously growing volumes by a variety of data sources and a variety of application fields rely on rapid analysis of such data. Existing systems such as PostGIS or MobilityDB usually build on relational database systems, thus, inheriting their scale-out characteristics. As a consequence, big spatiotemporal data scenarios still have limited support even though many query types can easily be parallelized. In this paper, we propose our vision of a native serverless data processing approach for spatiotemporal data: We break down queries into small subqueries which then leverage the near-instant scaling of Function-as-a-Service platforms to execute them in parallel. With this, we partially solve the scalability needs of big spatiotemporal data processing.
@@ 1. Introduction
Nowadays, spatiotemporal data are being produced in exponentially growing volumes in a variety of application scenarios. Examples of spatiotemporal data include mobility traces of cars, cyclists, pedestrians, planes, or ships in transport, movement patterns of wildlife in animal observation, Earth observation data, or satellite movement data (Liang et al., 2024). Aside from data values such as a temperature measurement or a boolean “something has a value here,” spatiotemporal data contain additional space and time attributes (Alam et al., 2022). Here, “space attributes” refers to locations usually indicated by latitude and longitude values and sometimes an additional geofence indicating a particular area for 2D spatial data or its 3D variant with an additional elevation level or 3D shapes. Time attributes, in contrast, provide a temporal context, usually based on timestamps but possibly also based on timespans. Such spatiotemporal data are usually encoded as either trajectory or raster data and stored in respective database systems. In this paper, we focus on trajectory data.
While there are different ways for processing these data, e.g., in batch mode, many application domains rely on interactive analysis of big spatiotemporal data with queries that filter and aggregate by space and time attributes at the same time. Existing solutions for this build on PostGIS or MobilityDB and, thus, inherit the limited scalability of the underlying relational PostgreSQL (Rese et al., 2025). Recent work by Bakli et al. (Bakli et al., 2025) here represents the state-of-the-art. Through complex sharding mechanisms and careful query planning, they managed to scale MobilityDB to up to eight machines – an impressive feat.
@@
At the same time, many query types can easily be broken down into vast numbers of subqueries operating on small data shards only, e.g., based on geographic regions or other spatial or temporal attributes, thus, parallelizing the query execution. Today, the go-to solution for these “embarrassingly parallel” (Hellerstein et al., 2019) tasks is Function as a Service (FaaS) in which developers write simple, stateless functions whose operation including on-demand scaling is then fully managed by the cloud provider (Bermbach et al., 2021).
In this paper, we hence sketch out our vision of a serverless data processing approach for spatiotemporal big data analytics, thus, leverage the rapid on-demand scaling of FaaS function instances for parallelizing query processing. We specifically do not make assumptions about how the data is structured, or whether it is being altered at some point during the analysis. We address the challenges faced in data analytics with complex and read-heavy workloads where the sheer amount of data that needs to be analyzed is way beyond anything a single node can manage efficiently. For this, we propose to break down queries into subqueries following the MapReduce model (Dean and Ghemawat, 2008), automatically deploy these on a FaaS platform, and merge the set of subresults into a single response (see Figure 1). In our proposed approach, a dedicated coordinator function analyzes the query and then breaks it down based on spatial or temporal characteristics or a combination of both.
This idea is not completely new, in fact, our work was partially inspired by similar approaches for serverless processing of tabular data (Lounissi et al., 2025; Bodner et al., 2025). To our knowledge, it has not yet been proposed for spatiotemporal data – a data type which needs to be dedicated sharding strategies which can consider spatial and temporal characteristics. Existing sharding approaches are not directly applicable. The reason for this is that storing spatial data already means reducing 2D or 3D data to a 1D sequence of values, which also makes indexing hard (Rese et al., 2025), while the time context adds at least one more dimension. Processing spatiotemporal big data on a single node, however, leads to slow response times as it stores and processes all data alone leading to nodes becoming slow which impacts any other workloads.
@@ 2. Approach
We propose to use a shared storage as data layer and shift query processing to FaaS applications. Striving for a modular design which allows for later integration of our approach into diverse architectures, i.e., independent of the implemented execution framework, we use FaaS for processing and blob storage for storing results. Our architecture design, depicted in Figure 2, is as follows.
A spatiotemporal query initiated by the client, e.g., “Identify the flight-heaviest regions in the last six months” is our starting point. A starter function provides an HTTP endpoint to receive the query and invokes the centralized coordinator. The coordinator splits the query into subqueries and initializes the worker functions. Queries such as SELECT can easily be distributed and executed on data shards, while others, e.g., ORDER BY or AVG, depend on context, i.e., they require the complete result of the previous subquery to return correct result and therefore can only be executed by single node. This is why we design our architecture in a way that it supports all kinds of queries, parallelizable and non-parallelizable. For that, the coordinator manages waves of worker functions for each subquery with the amount of worker functions per wave dependent on the parallelizability of the subquery.
@@ 3. Conclusion & Future Work
In this paper, we proposed our vision of a native serverless data processing approach for spatiotemporal data analysis with complex and read-heavy workloads: The proposed approach breaks down queries into small subqueries which then leverage the near-instant scaling of Function-as-a-Service platforms for parallel execution.